OpenStudy (anonymous):

dang sorry looked at wrong question here is real question what is perpendicular to -2/3x+7 and passes through (2,1)

OpenStudy (anonymous):

perpendicular slopes are reciprocals with a change of sign i.e. if the line in question is y = 2x a perpendicular line would be y = -1/2x then plug in the point (2,1) into the equation y = mx + b and you find b

OpenStudy (anonymous):

when i plugged it in i got 1=3 wat do i do after?

OpenStudy (anonymous):

what did you get for the perpendicular slope m?

OpenStudy (anonymous):

you have one unknown, b. after you have figured the slope, you have x and y so plug in to y = mx + b. solve for b.

OpenStudy (anonymous):

my slope was 3/2x

OpenStudy (anonymous):

ok thats correct. plug in (2,1) so you should have \[y = mx + b\] and if you plug in correctly it will become \[1 = (3/2)*2 + b\] solve for b. then go back to \[y = mx + b\] use that b to plug in and also the slope you just found for a general equation of the line.